When the repeating decimal 0.4848 ... is written in simplest fractional form,
what is the sum of the numerator and denominator?
A. 32; B. 49; C. 137; D. 148; E. none

Answer by Alan3354(21605)

When the repeating decimal 0.4848 ... is written in simplest fractional form,
what is the sum of the numerator and denominator?
-----------
n = 0.4848484848
100n= 48.484848484
- 1n = - 0.48484848
------------------ Subtract
99n = 48
n = 48/99 = 16/33
--------
B. 49